Circular No. 7475

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000db IN NGC 3949 S. Nakano, Sumoto, Japan, reports the discovery by M. Aoki, Tsukioka,

Toyama, of an apparent supernova (mag 14.3) on a CCD image taken Aug. 6.498 UT and confirmed on six unfiltered CCD frames taken Aug. 7.444 (limiting mag 18.5), using an 0.43-m f/6 reflector + CCD and an automated supernova survey system developed by Aoki. The supernova is located at α = 11^{h}53^{m}39^{s}.75, δ = +47^{o}51′46′′.6 (equinox 2000), which is 20′′ northwest of the center of NGC 3949. No star is present at this location on Aoki’s 19 frames taken between 1996 and May 2000.

COMET C/1999 S4 (LINEAR) D. Schleicher and L. Woodney, Lowell Observatory, write: “We used

the Hall 1.1-m telescope (+ 95′′-diameter aperture) at Lowell Observatory to obtain narrowband photometry of comet C/1999 S4 (centered at the nominal ephemeris position; cf. MPEC 2000-O02), yielding the following production rates: log Q(OH) = 27.4; log Q(CN) = 24.4; log Afρ = 1.5 (cf. IAUC 7342). The equivalent log Q (water; vectorial) is 27.6. The values are 8–10 times less than those measured on July 13 (cf. IAUC 7455).”

Though necessarily uncertain, the following ephemeris extension to IAUC 7455 (from elements on MPC 40988) is given as a finding aid.

2000 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β
Aug. 14 12^{h}35.^{m}59 − 7^{o}21.^{′}5 1.009 0.848 49.^{o}6 65.^{o}5
16 12 37.17 − 8 48.8 1.081 0.866 48.7 61.5
18 12 38.48 −10 04.6 1.152 0.884 47.7 57.8
20 12 39.57 −11 11.0 1.222 0.904 46.5 54.4
22 12 40.49 −12 09.8 1.291 0.924 45.3 51.1
24 12 41.30 −13 02.4 1.359 0.945 44.1 48.0
28 12 42.64 −14 32.8 1.490 0.991 41.4 42.4
Sept. 1 12 43.76 −15 48.7 1.615 1.039 38.6 37.3
5 12 44.76 −16 54.2 1.733 1.088 35.8 32.8
9 12 45.70 −17 52.2 1.845 1.140 32.9 28.7
13 12 46.61 −18 44.6 1.951 1.193 30.1 25.1
17 12 47.52 −19 32.9 2.049 1.247 27.4 21.8
21 12 48.42 −20 18.1 2.141 1.301 24.9 18.9
25 12 49.33 −21 01.1 2.227 1.356 22.5 16.4
29 12 50.24 −21 42.5 2.306 1.411 20.4 14.3
Oct. 3 12 51.14 −22 22.6 2.379 1.467 18.7 12.6

2000 August 7 \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT Elizabeth O. Waagen